If you traveling at 80mph how long would it take you to go 50 miles
Answer by oberobic(2304)

d = r*t
50 = 80*t
t = 50/80 = 5/8 hr
.
5/8 hr * 60 min/hr = 300/8 = 150/4 = 75/2 = 37 1/2 minutes
.
Done.